OpenStudy (anonymous):

The dimensions of four different cylinders are given below. Which cylinder has the greatest volume? Cylinder 1: r = 4ft, h = 3ft Cylinder 2: r = 2ft, h = 5ft Cylinder 3: r = 3ft, h = 4ft Cylinder 4: r = 5ft, h = 1ft A)Cylinder 1 B)Cylinder 2 C)Cylinder 3 D)Cylinder 4

OpenStudy (anonymous):

i give medal to the one with the best answer

Parth (parthkohli):

One with the most \(\rm radius \times radius \times height\)

OpenStudy (anonymous):

which one is that they have the same xradius and everything

Parth (parthkohli):

Rephrase please?

OpenStudy (anonymous):

like is it a ,b ,c or d

OpenStudy (anonymous):

volume=pi*r*r*h option A is the required answer

Parth (parthkohli):

I dunno.

OpenStudy (anonymous):

thank you guys for your help

Parth (parthkohli):

No, we didn't help ye.